OpenStudy (anonymous):

what is equivalent to 2(x-5)+4x-18=4x

OpenStudy (anonymous):

help plzz

OpenStudy (anonymous):

are we sloving for x? if so x= 14

OpenStudy (anonymous):

i need to find what is equivalent to it

OpenStudy (anonymous):

so then i am assuming they just want you to make the problem smaller so basically all your doing is getting rid of the "(" ")" 2(x-5+4x-18) = 4x 2x-10+8x-36 =4x *then add like terms* 10x-46 =4x <--- so i believe that would be your answer cause it is equivalent to your original problem.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

You're Welcome(: if you need any help with math let me know ill do my best to help(: